40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Sep 16 19:48:52 2016, Lydia, Update, General, Alignment status 
    Reply  Sat Sep 17 02:00:23 2016, rana, Update, SUS, Alignment status 
       Reply  Sat Sep 17 16:51:01 2016, Lydia, Update, SUS, Alignment status Screenshot_from_2016-09-17_16-45-00.png
          Reply  Sun Sep 18 16:18:05 2016, rana, Update, SUS, Alignment status 
             Reply  Mon Sep 19 11:11:43 2016, ericq, Update, SUS, Alignment status 
                Reply  Mon Sep 19 22:03:10 2016, ericq, Update, General, Further recovery progress 
                   Reply  Tue Sep 20 10:45:06 2016, rana, Update, General, Further recovery progress 
                      Reply  Wed Sep 21 01:08:02 2016, ericq, Update, General, Further recovery progress 
                         Reply  Wed Sep 21 09:04:57 2016, Steve, Update, General, 8 hours recovery progress 8hrs.png
                            Reply  Thu Sep 22 15:44:21 2016, Steve, Update, General, ITMX magnets are stucked again StuckAgain.png
                               Reply  Thu Sep 22 20:01:47 2016, Lydia, Update, General, ITMX freed, all optics kicked 
                                  Reply  Thu Sep 22 22:52:08 2016, ericq, Update, General, Damping found to be on 
Message ID: 12502     Entry time: Sat Sep 17 16:51:01 2016     In reply to: 12501     Reply to this: 12503  
Author: Lydia 
Type: Update 
Category: SUS 
Subject: Alignment status 

Here's the timeseries plots. I've zoomed in to right after the problem- did you want before? We pretty much know what happened: c1susaux was restarted from the crate but the damping was on, so as soon as the machine came back online the damping loops sent a huge signal to the coils. (Also, it seems to be down again. Now we know what to do first before keying the crate.) It seems like both right side magnets are stuck, and this could probably be fixed by moving the yaw slider. Steve advised that we wait for an experienced hand to do so. 

Quote:

All is not lost. I've stuck and unstuck optics around a half dozen times. Can you please post the zoomed in time series (not trend) from around the time it got stuck? Sometimes the bias sliders have to be toggles to make the bias correct. From the OSEM trend it seems like it got a large Yaw bias. May also try to reseat the satellite box cables and the cable from the coil driver to the cable breakout board in the back of the rack.

 

Attachment 1: Screenshot_from_2016-09-17_16-45-00.png  28 kB  | Hide | Hide all
Screenshot_from_2016-09-17_16-45-00.png
ELOG V3.1.3-